OpenStudy (anonymous):

The gardener mixed up the petunias he bought at the plant shop. None of them are blooming, but she knows she bought 3 purple, 15 red, and 13 white. if he just selects one at random, what is the probability that it is red? explain.

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

i think it should be \(\Large P(\color{red}{red})=\dfrac{no~of~\color{red}{red}~flowers}{\tt \text{total number of flowers}}\)

OpenStudy (anonymous):

15/28 ???

Directrix (directrix):

Check the total number of flower plants again.

OpenStudy (anonymous):

yeah i tried doing it in my head lol

OpenStudy (anonymous):

and i didnt add the 3

Directrix (directrix):

>>> 15/28 ??? So, what should that be?

OpenStudy (anonymous):

its a really big decimal.....

Directrix (directrix):

As a fraction first.

OpenStudy (anonymous):

0.535714286

Directrix (directrix):

15 over what?

OpenStudy (anonymous):

is what i got

OpenStudy (anonymous):

15/28

Directrix (directrix):

28 is not correct for total flowers

OpenStudy (anonymous):

well mathmath told me to add all of the numbers

OpenStudy (anonymous):

so thats how i got a denominator of 28

Directrix (directrix):

Yes. I think there is an error in the sum: 3 + 15 + 13 = what

OpenStudy (anonymous):

31

OpenStudy (mathmath333):

so it should be \(P(red)=15/31\)

OpenStudy (anonymous):

so is it 15/31 or not??

OpenStudy (mathmath333):

lol it is